OpenStudy (togundeji):

9. In ΔABC, m∠A = 4x, m∠B = 3x + 1, m∠C = 7x + 11. List the sides of ΔABC in order from shortest to longest. AB;BC;AC AC;BC;AB BC;AB;AC AC;AB;BC

OpenStudy (togundeji):

WILL GIVE BEST ANSWER A MEDAL

563blackghost (563blackghost):

We know that the total degree of a triangle is 180 so in order to find x we would add up all the angles and simplify....\[(4x) +(3x+1)+(7x+11)=180\]

OpenStudy (togundeji):

That's what I did but I've got something that is not one of the options

563blackghost (563blackghost):

Well I'll do it with you ^^.... So we would first combine like terms....so...\[4x+3x+7x+1+11=180\]which would simplify too...\[14x+12=180\]We then subtract 12 from each side to get 14x by itself.... \[14x=168\]We then divide by 14...so what would x be?

OpenStudy (togundeji):

0.08333333333

563blackghost (563blackghost):

Incorrect.... it would actually be 12... since 14 times 12 equals 168... Now we would input 12 where x is....and then we would find the smallest angle as well as the largest...

OpenStudy (togundeji):

I thank you for the correction I just need to find the small angles and the biggest angles.

563blackghost (563blackghost):

So we input.... \[4(12)\] \[3(12)+1\] \[7(12)+11\] Now we simplify...

OpenStudy (amity):

<A+<B+<C=180, 4x+3x+17x+11=180,,,,,x=12 hence angles A,B C are 48,37,,95 in degres respectively

OpenStudy (togundeji):

so whats the answer, amity?

563blackghost (563blackghost):

Its best to understand than given an answer...

563blackghost (563blackghost):

im sorry but I must go...

563blackghost (563blackghost):

And @amity please dont give out the answer it would be in violation of the rules...

OpenStudy (amity):

Sides opposite to the greater angle are greater in triangle by property of inequality hence answer will be AC,BC,AB....2nd option is correct

OpenStudy (amity):

Sides opposite to the greater angle are greater in triangle by property of inequality hence answer will be AC,BC,AB....2nd option is correct

OpenStudy (togundeji):

I am passing geometry now with a B!! I already turned in my test but I'm sure it will some how go down to a D. @amity @563blackghost